OpenStudy (anonymous):

An asteroid follows a hyperbolic path defined by the equation 16x^2 − 9y^2 = 576. If one of its foci is the sun’s position, the minimum distance between the asteroid and the sun is(how many?) AU. (Note that x and y are expressed in astronomical units, AU.)

OpenStudy (triciaal):

not a pro with these related to a 3-4-5 triangle do you happen to have choices?

OpenStudy (anonymous):

No I don't :(

OpenStudy (anonymous):

@iambatman can you help please?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@TuringTest @wio

OpenStudy (triciaal):

@jim_thompson5910 @Directrix @PRAETORIAN.10

ganeshie8 (ganeshie8):

As a start : change the given equation into standard form \[\large \dfrac{x^2}{a^2} - \dfrac{y^2}{b^2}=1\]

ganeshie8 (ganeshie8):

divide both sides by 576 so that the right hand side becomes 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Jhannybean

ganeshie8 (ganeshie8):

\[16x^2 − 9y^2 = 576\] dividing \(576\) through out you get \[\dfrac{16x^2}{576} − \dfrac{9y^2}{576} = \dfrac{576}{576}\]

ganeshie8 (ganeshie8):

which is same as \[\dfrac{x^2}{36} - \dfrac{y^2}{64} = 1\]

ganeshie8 (ganeshie8):

compare this with the standard form and see if you can eyeball the values of \(a\) and \(b\)

OpenStudy (anonymous):

6 & 8?

ganeshie8 (ganeshie8):

Yes! a = 6 b = 8 so the x coordinate of vertex is 6

ganeshie8 (ganeshie8):

Next find the x coordinate of focus by using the relation : \[\large c^2 = a^2+b^2\]

OpenStudy (anonymous):

100

ganeshie8 (ganeshie8):

\[\large c^2 = a^2+b^2\] \[\large c^2 = 6^2+8^2\] \[\large c^2 = 100\] \[\large c = ?\]

OpenStudy (anonymous):

10

ganeshie8 (ganeshie8):

correct!

ganeshie8 (ganeshie8):